Grade 12th passAlgebra

Find five rational numbers between.
¼ and ½ …..........

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Let us make the denominators same, say 24.
i.e., 1⁄4 and 1⁄2 can be written as:
1⁄4= (1×6)/(4×6)=6/24
1⁄2=(1×12)/(2×12)=12/24
Five rational numbers between 1⁄4 and 1⁄2 = five rational numbers between 6/24 and 12/24 Therefore, Five rational numbers between 6/24 and 12/24 = 7/24, 8/24, 9/24, 10/24, 11/24

Thanks